# Laplace measure transitions and ghosts for meromorphic functions

### 05 Apr 2026 v1

@@ 6 Ghost transitions
The examples in section 5 illustrate two types of behavior. For 5.2, 5.3 and 5.4, the presence of poles over the separatrix accounted for the difference between $\mu_{r}$ and $\mu_{l}$ and the transition formula was valid in the appropriate sense. The second type of behavior entailed the loss of Laplace representation upon crossing a separatrix containing no poles: example 5.1 of the $\zeta$ function on the strip $\Omega_{0,1}$ and $\zeta(s)/s$ on $\Omega_{-1,0}.$ This leads to the following question:
@@
$\gamma_{M}=\bigg\{\zeta\in\mathbb{C}:|\sigma|=\frac{\pi}{2\tau},\tau\geq M\bigg\}\cup\bigg\{\zeta\in\mathbb{C}:\tau=M,|\sigma|\leq\frac{\pi}{2M}\bigg\}.$ (31)
@@
$I(s)=\frac{1}{2\pi i}\int_{\gamma_{M}}\frac{e^{e^{-\zeta^{2}}}}{s-\zeta}\,d\zeta$ (32)
@@
Denote by $s_{1}^{\pm}=\pm\frac{\pi}{2y_{1}}+iy_{1}$ the points on $\gamma_{M}$ closest to $s=iy.$ By considering the isosceles triangle with vertices at $s$ and $s_{1}^{\pm}$ it follows that $d\geq\frac{\pi}{2y_{1}},$ where $d=\inf_{\zeta\in\gamma_{M}}|\zeta-iy|.$ Since $\lim_{y\to+\infty}|y-y_{1}|=0$ uniformly, there is a positive constant $C_{1}$ such that $C_{1}y\geq y_{1}$ for all pairs $(y,y_{1}).$ It follows that for all $\zeta\in\gamma_{M}$
@@
For $\Re(s)>\frac{\pi}{2M}$ we use the Laplace representation of the Cauchy kernel
@@
and express $F(s)$ by
$F(s)=\frac{1}{2\pi i}\int_{\gamma_{M}}e^{e^{-\zeta^{2}}}\int_{0}^{+\infty}e^{-st}e^{\zeta t}\,dt\,d\zeta.$ (35)
@@
$F(s)=\int_{0}^{+\infty}e^{-st}\Bigg(\frac{1}{2\pi i}\int_{\gamma_{M}}e^{e^{-\zeta^{2}}}e^{\zeta t}\,d\zeta\Bigg)\,dt,\quad\Re(s)>\frac{\pi}{2M}.$ (36)
@@
where $\partial_{\pi}$ denotes the right half-strip delimited by $\Re(s)=0$ and $|\Im(s)|<\pi,$ $\mathcal{H}$ is the classical Hankel contour for the $\Gamma$ function ( see section 13.2.4. of Krantz [20]) and $\gamma$ is the Euler-Mascheroni constant. Therefore $f(t)\equiv f_{r}(t)-f_{l}(t)$ is not identically zero and is non-trivial on a set of positive Lebesgue measure. This concludes the proof.
@@ 6.1 Connection to the heat equation
The function given by
$H:\mathbb{R}\times[0,+\infty)\longrightarrow\mathbb{C}$ (40)
@@
Interchanging the order of integration yields
$H(x,t)=\frac{1}{2\pi i}\int_{\gamma_{M}}e^{e^{-\zeta^{2}}}\int_{-\infty}^{+\infty}e^{ix\lambda}e^{-t\lambda^{2}+\zeta\lambda}\,d\lambda\,d\zeta=\frac{1}{\sqrt{4\pi t}i}\int_{\gamma_{M}}e^{e^{-\zeta^{2}}}e^{-\frac{(x-i\zeta)^{2}}{4t}}\,d\zeta$ (42)
Finally, performing a change of variable $i\zeta=\omega$ in (42) we recover the known non-trivial solution to the heat IVP (37)
@@
which was originally constructed by Chung in [8]. ∎

### 07 May 2026 v2

@@ 6 Ghost transitions
The examples in section 5 illustrate two types of behavior. For 5.2, 5.3 and 5.4, the presence of poles on the separatrix accounted for the difference between $\mu_{r}$ and $\mu_{l}$ and the transition formula was valid in the appropriate sense. The second type of behavior entailed the loss of Laplace representation upon crossing a separatrix containing no poles: this is the case of example 5.1, dealing with the $\zeta$ function on the strip $\Omega_{0,1}$ and $\zeta(s)/s$ on $\Omega_{-1,0}.$ This leads to the following question:
@@
$\gamma_{M}=\bigg\{\zeta\in\mathbb{C}:|\sigma|=\frac{\pi}{2\tau},\tau\geq M\bigg\}\cup\bigg\{\zeta\in\mathbb{C}:\tau=M,|\sigma|\leq\frac{\pi}{2M}\bigg\}.$
@@
$I(s)=\frac{1}{2\pi i}\int_{\gamma_{M}}\frac{e^{e^{-\zeta^{2}}}}{s-\zeta}\,d\zeta$
@@
Denote by $s_{1}^{\pm}=\pm\frac{\pi}{2y_{1}}+iy_{1}$ the points on $\gamma_{M}$ closest to $s=iy.$ By considering the isosceles triangle with vertices at $s$ and $s_{1}^{\pm}$ it follows that $d\geq\frac{\pi}{2y_{1}},$ where $d=\inf_{\zeta\in\gamma_{M}}|\zeta-iy|.$ Since $\lim_{y\to+\infty}|y-y_{1}|=0$ uniformly, there is a positive constant $C_{1}$ such that $C_{1}y\geq y_{1}$ for all pairs $(y,y_{1}).$ It follows that, for all $\zeta\in\gamma_{M}$,
@@
We use the Laplace representation of the Cauchy kernel
@@
and, for $\Re(s)>\frac{\pi}{2M}$, express $F(s)$ by
$F(s)=\frac{1}{2\pi i}\int_{\gamma_{M}}e^{e^{-\zeta^{2}}}\int_{0}^{+\infty}e^{-st}e^{\zeta t}\,dt\,d\zeta.$
@@
$F(s)=\int_{0}^{+\infty}e^{-st}\Bigg(\frac{1}{2\pi i}\int_{\gamma_{M}}e^{e^{-\zeta^{2}}}e^{\zeta t}\,d\zeta\Bigg)\,dt,\quad\Re(s)>\frac{\pi}{2M}.$
@@
where $\partial_{\pi}$ denotes the right half-strip delimited by $\Re(s)=0$ and $|\Im(s)|<\pi,$ $\mathcal{H}$ is the classical Hankel contour for the $\Gamma$ function (see section 13.2.4. of Krantz [20]) and $\gamma$ is the Euler-Mascheroni constant. Therefore $f(t)\equiv f_{r}(t)-f_{l}(t)$ is not identically zero and is non-zero on a set of positive Lebesgue measure. This concludes the proof.
@@ 6.1 Connection to the heat equation
The function $H:\mathbb{R}\times[0,+\infty)\to\mathbb{C}$ given by
@@
Interchanging the order of integration in (32) yields
$H(x,t)=\frac{1}{2\pi i}\int_{\gamma_{M}}e^{e^{-\zeta^{2}}}\int_{-\infty}^{+\infty}e^{ix\lambda}e^{-t\lambda^{2}+\zeta\lambda}\,d\lambda\,d\zeta=\frac{1}{\sqrt{4\pi t}i}\int_{\gamma_{M}}e^{e^{-\zeta^{2}}}e^{-\frac{(x-i\zeta)^{2}}{4t}}\,d\zeta.$
Performing the change of variable $i\zeta=\omega$ we recover the known non-trivial solution to the heat IVP (29)
@@
originally constructed by Chung in [8]. ∎